OpenStudy (vauthriel):

x^radlogx = 10^6; solve for x

OpenStudy (vauthriel):

\[x ^{\sqrt{logx} } = 10^6\]

OpenStudy (asapolous):

1000000000000

OpenStudy (pythagoras123):

take log on both sides \[\sqrt {lgx} \times lgx = 6\] Let lgx = u \[u^\frac{ 3 }{ 2} = 6\] \[u=6^\frac{ 2 }{ 3}\] \[x=10^{6^{\frac{ 2 }{ 3}}}\] \[x=10000\]